# Hybrid estimation for a mixed fractional Black-Scholes model with random effects from discrete time observations

### v3

@@ 7.1 Proof of Lemma 1
$\rho(k)=\mbox{Cov}(\Delta M_{k}^{H,i},\Delta M_{0}^{H,i})=\sigma^{2}h\mathbf{1}_{\{k=0\}}+\gamma^{2}\rho_{H}(k),$ (7.1)
where $\rho_{H}(k)=\mathbb{E}[\Delta B^{H,i}_{k}\Delta B^{H,i}_{0}]=\frac{1}{2}\left[|k+1|^{2H}+|k-1|^{2H}-2|k|^{2H}\right]$. In the case $H\in\left(\frac{1}{2},1\right)$, the asymptotic behavior of $\rho_{H}(k)$ is as follows
@@ 7.2 Proof of Proposition 1
$\frac{1}{n}\sum_{k=0}^{n-1}\Delta Y_{k}=\theta h+\frac{1}{n}\sum_{k=0}^{n-1}M^{H,i}_{k}\xrightarrow{a.s.}\theta h,\quad\mathrm{as}\quad n\to\infty.$
@@ 7.4 Proof of Proposition 3
Asymptotic normality of $U^{i}_{n}:=\left(\hat{\theta}_{i}^{2},\,\xi_{n}^{i},\eta_{n}^{i},\zeta^{i}_{n}\right)^{\intercal}$
@@
$\hat{\theta}_{i}^{2}-\theta_{i}^{2}$ $=L_{n,\theta}+Q_{n,\theta}$ $:=\dfrac{2\theta h}{n}\sum\limits_{k=0}^{n-1}\Delta M_{k}^{H,i}+\Big(\dfrac{1}{nh}\sum_{k=0}^{n-1}\Delta M_{k}^{H,i}\Big)^{2},$
@@
where $\Sigma:=\mathrm{Cov}\big(U^{i}_{\infty}\big)=\mathrm{Var}\big(\theta_{i}^{2}\big)\,\mathcal{H}\,\mathcal{H}^{\top}$, with $\mathcal{H}:=\big(1,h^{2},h^{2},4h^{2}\big)^{\top}$.
@@ 7.5 Proof of Proposition 4
Now, since the quantities $V_{\infty}$, $\xi_{\infty}$, $\eta_{\infty}$ and $\zeta_{\infty}$ depend on the unknown expectation $\mathbb{E}\big(\theta_{i}^{2}\big)$, we replace them with their averaged statistics $\bar{V}_{n,N},\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N}$, which are based on the observed subjects. Therefore, we obtain the estimators $\hat{H}$, $\hat{\gamma}^{2}$ and $\sigma^{2}$ given by (3.8), (3.9) and (3.10).
@@ 7.6 Proof of Theorem 1
with $\Sigma=\mathrm{Var}(\theta_{1}^{2})\,\mathcal{H}\,\mathcal{H}^{\top}$. Moreover, as it was shown in the previous proof, the function $g$ is well defined at $U_{\infty}$ since $H>\frac{1}{2}$ and $\gamma>0$. Then, by the same arguments, we deduce that $g$ is continuously differentiable in a neighborhood of $U_{\infty}$. Therefore the multivariate delta method yields
@@
$\sqrt{N}\big(\widehat{\Theta}-\Theta\big)=\sqrt{N}\big(g(\bar{U}_{n,N})-g(U_{\infty})\big)\xrightarrow{d}\mathcal{N}\big(0,\,J\Sigma J^{\top}\big),$
where $J=Dg$ is the Jacobian of $g$ evaluated at $U_{\infty}$, i.e
@@
$J=\begin{pmatrix}\frac{\partial\hat{H}}{\partial\bar{\xi}_{n,N}}&\frac{\partial\hat{H}}{\partial\bar{\eta}_{n,N}}&\frac{\partial\hat{H}}{\partial\bar{\zeta}_{n,N}}&\frac{\partial\hat{H}}{\partial\bar{V}_{n,N}}\\ \frac{\partial\hat{\gamma}^{2}}{\partial\bar{\xi}_{n,N}}&\frac{\partial\hat{\gamma}^{2}}{\partial\bar{\eta}_{n,N}}&\frac{\partial\hat{\gamma}^{2}}{\partial\bar{\zeta}_{n,N}}&\frac{\partial\hat{\gamma}^{2}}{\partial\bar{V}_{n,N}}\\ \frac{\partial\hat{\sigma}^{2}}{\partial\bar{\xi}_{n,N}}&\frac{\partial\hat{\sigma}^{2}}{\partial\bar{\eta}_{n,N}}&\frac{\partial\hat{\sigma}^{2}}{\partial\bar{\zeta}_{n,N}}&\frac{\partial\hat{\sigma}^{2}}{\partial\bar{V}_{n,N}}\end{pmatrix}_{|{U_{\infty}}}.$
@@
In order to simplify notation, we denote $\bar{A}:=\bar{\zeta}_{n,N}-4h^{2}\bar{V}_{n,N},$ and $\bar{B}:=\bar{\eta}_{n,N}-h^{2}\bar{V}_{n,N}.$ Then, the estimator $\hat{H}$ can be expressed by $\widehat{H}=\frac{1}{2\ln 2}\big(\ln\bar{A}-\ln\bar{B}\big)$ and its derivatives with respect to $\bar{U}_{n,N}$ are given by
$\frac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}=\frac{h^{2}}{2\ln 2}\!\left(\frac{1}{\bar{B}}-\frac{4}{\bar{A}}\right),\ \ \frac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}=0,\ \ \frac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}=-\frac{1}{2\ln 2}\,\frac{1}{\bar{B}},\ \ \frac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}=\frac{1}{2\ln 2}\,\frac{1}{\bar{A}}.$
@@
Evaluating at $U_{\infty}$ and using $A=2^{\,2H}B$ and $B=\gamma^{2}D_{H}$ with $D_{H}:=h^{2H}\big(2^{2H-1}-1\big)$, yields
$\frac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}=\frac{h^{2}}{2\ln 2}\,\frac{2^{2H}-4}{2^{2H}\gamma^{2}D_{H}},\quad\frac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}=0,\quad\frac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}=-\frac{1}{2\ln 2}\,\frac{1}{B},\quad\frac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}=\frac{1}{2\ln 2}\,\frac{1}{2^{\,2H}B}.$
@@
Next, for the estimator $\widehat{\gamma}^{\,2}=\frac{\bar{B}}{D_{\hat{H}}}$, the chain rule gives for any component $u\in\{\bar{V}_{n,N},\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N}\}$,
$\frac{\partial\widehat{\gamma}^{\,2}}{\partial u}=\frac{\partial_{u}\bar{B}}{D_{\widehat{H}}}-\bar{B}\,\frac{D^{\prime}_{\widehat{H}}}{D_{\widehat{H}}^{2}}\,\frac{\partial\widehat{H}}{\partial u},\qquad D^{\prime}_{H}:=\frac{\partial D_{H}}{\partial H}=h^{2H}\!\Big[2\ln(h)\big(2^{2H-1}-1\big)+(2\ln 2)\,2^{2H-1}\Big].$
@@
Using $\frac{\partial\bar{B}}{\partial\bar{V}_{n,N}}=-h^{2}$, $\partial_{\bar{\eta}}\bar{B}=1$, and $\frac{\partial\bar{B}}{\partial\bar{\xi}_{n,N}}=\frac{\partial\bar{B}}{\partial\bar{\zeta}_{n,N}}=0$, and evaluating at $U_{\infty}$ with $B=\gamma^{2}D_{H}$ and $A=2^{\,2H}B$, we obtain
$\frac{\partial\widehat{\gamma}^{\,2}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}$ $=-\,\frac{h^{2}}{D_{H}}+\frac{h^{2}}{2\ln 2}\!\left(\frac{4}{2^{\,2H}}-1\right)\frac{D^{\prime}_{H}}{D_{H}^{2}},\quad\frac{\partial\widehat{\gamma}^{\,2}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}=0,$
@@
$\frac{\partial\widehat{\gamma}^{\,2}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}$ $=\frac{1}{D_{H}}+\frac{D^{\prime}_{H}}{2\ln 2}\,\frac{1}{D_{H}^{2}},\quad\frac{\partial\widehat{\gamma}^{\,2}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}=-\,\frac{D^{\prime}_{H}}{2\ln 2}\,\frac{1}{2^{\,2H}D_{H}^{2}}.$
Finally, for $\widehat{\sigma}^{\,2}=\frac{\bar{\xi}-h^{2}\bar{V}}{h}-\frac{\widehat{\gamma}^{\,2}h^{2\widehat{H}}}{h},$ we use
$\frac{\partial}{\partial u}\!\big(\widehat{\gamma}^{\,2}h^{2\widehat{H}}\big)=\big(\frac{\partial}{\partial u}\widehat{\gamma}^{\,2}\big)h^{2\widehat{H}}+\widehat{\gamma}^{\,2}h^{2\widehat{H}}(2\ln h)\,\frac{\partial\widehat{H}}{\partial u}.$
@@
Evaluating at $U_{\infty}$, which means that $h^{2\widehat{H}}\to h^{2H}$ and $\gamma^{2}h^{2H}=\frac{B}{D_{H}}$, we obtain the following
@@
$\frac{\partial\widehat{\sigma}^{\,2}}{\partial V_{n,N}}\Big|_{U_{\infty}}$ $=-\,h\;+\;h^{\,2H-1}\!\Bigg[\frac{h^{2}}{D_{H}}-\frac{h^{2}}{2\ln 2}\!\left(1-\frac{4}{2^{\,2H}}\right)\left(\!-\frac{D^{\prime}_{H}}{D_{H}^{2}}+\frac{2\ln(h)}{D_{H}}\!\right)\Bigg],$
@@
$\frac{\partial\widehat{\sigma}^{\,2}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}$ $=-\,h^{\,2H-1}\,\frac{1}{D_{H}}\left[\,1-\frac{2\ln(h)}{2\ln 2}+\frac{D^{\prime}_{H}}{2\ln 2}\,\frac{1}{D_{H}}\right],$
$\frac{\partial\widehat{\sigma}^{\,2}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}$ $=\;h^{\,2H-1}2^{-2H}\!\left[\,\frac{D^{\prime}_{H}}{2\ln 2}\,\frac{1}{D_{H}^{2}}-\frac{2\ln(h)}{2\ln 2}\,\frac{1}{D_{H}}\right].$

### v4

@@ 7.1 Proof of Lemma 1
$\rho(k)=\mathrm{Cov}(\Delta M^{H,i}_{k},\Delta M^{H,i}_{0})=\begin{cases}\sigma^{2}h+\gamma^{2}h^{2H}&\text{si }k=0,\\ \gamma^{2}\rho_{H}(k)&\text{si }k\neq 0,\end{cases}$ (7.1)
where $\rho_{H}(k)=\frac{1}{2}\left(|k+1|^{2H}+|k-1|^{2H}-2|k|^{2H}\right)$ for $k\geq 1$. In the case $H\in\left(\frac{1}{2},1\right)$, the asymptotic behavior of $\rho_{H}(k)$ is as follows
@@ 7.2 Proof of Proposition 1
$\frac{1}{n}\sum_{k=0}^{n-1}\Delta Y^{i}_{k}\xrightarrow{a.s.}\theta_{i}\,h,\quad\mathrm{as}\quad n\to\infty.$
@@ 7.4 Proof of Proposition 3
Asymptotic normality of $U^{i}_{n}:=\left(\widehat{\theta}_{i}^{2},\,\xi_{n}^{i},\eta_{n}^{i},\zeta^{i}_{n}\right)^{\intercal}$
@@
$\widehat{\theta}_{i}^{2}-\theta_{i}^{2}$ $=L_{n,\theta}+Q_{n,\theta}$ $:=\dfrac{2\theta h}{n}\sum\limits_{k=0}^{n-1}\Delta M_{k}^{H,i}+\Big(\dfrac{1}{nh}\sum_{k=0}^{n-1}\Delta M_{k}^{H,i}\Big)^{2},$
@@
Remark 5.
The normalization factor $n^{1-H}$ in the equation (7.7) (rather than the standard $\sqrt{n}$) arises because when $H>\frac{1}{2}$, the fBm component, which contributes errors of order $n^{H-1}$, dominates the Brownian motion component, which contributes errors of order $n^{-1/2}$. This different scaling is characteristic of statistical inference for fractional processes and distinguishes it from classical Brownian-based inference.
@@
where $\Sigma:=\mathrm{Cov}\big(U^{i}_{\infty}\big)=\mathrm{Var}\big(\theta_{i}^{2}\big)\,\mathcal{H}\,\mathcal{H}^{\top}$, with $\mathcal{H}:=\big(1,h^{2},h^{2},4h^{2}\big)^{\top}$. Finally, by combining (7.9) and (7.10), the Slutsky’s theorem A.1 yields
$\sqrt{N}\left(\bar{U}_{n,N}-U_{\infty}\right)\ \xrightarrow{d}\mathcal{N}(0,\Sigma),\quad\mathrm{as}\quad n\to\infty\ \ \mathrm{then}\ \ \ N\to\infty.$ (7.11)
@@ 7.5 Proof of Proposition 4
Now, since the quantities $V_{\infty}$, $\xi_{\infty}$, $\eta_{\infty}$ and $\zeta_{\infty}$ depend on the unknown expectation $\mathbb{E}\big(\theta_{i}^{2}\big)$, we replace them with their averaged statistics $\bar{V}_{n,N},\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N}$, which are based on the observed subjects. Therefore, we obtain the estimators $\widehat{H}$, $\widehat{\gamma}^{2}$ and $\sigma^{2}$ given by (3.9), (3.10) and (3.11).
@@ 7.6 Proof of Theorem 1
with $\Sigma=\mathrm{Var}(\theta_{1}^{2})\,\mathcal{H}\,\mathcal{H}^{\top}$. Moreover, as it was shown in the previous proof, the function $g$ is well defined at $U_{\infty}$ since $H>\frac{1}{2}$ and $\gamma>0$. Then, by the same arguments, we deduce that $g$ is continuously differentiable in a neighborhood of $U_{\infty}$. Therefore, the multivariate delta method yields
@@
$\sqrt{N}\bigl(\widehat{\Theta}-\Theta\bigr)=\sqrt{N}\bigl(g(\bar{U}_{n,N})-g(U_{\infty})\bigr)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\bigl(0,\,J\Sigma J^{\top}\bigr),$
where $J=Dg$ is the Jacobian of $g$ evaluated at $U_{\infty}$, i.e.,
@@
$J=\begin{pmatrix}\displaystyle\frac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}\\[12.0pt] \displaystyle\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}\\[12.0pt] \displaystyle\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}&\displaystyle\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}\end{pmatrix}$.
@@
In order to simplify notation, we denote $\bar{A}:=\bar{\zeta}_{n,N}-4h^{2}\bar{V}_{n,N}$ and $\bar{B}:=\bar{\eta}_{n,N}-h^{2}\bar{V}_{n,N}$. Then, the estimator $\widehat{H}$ can be expressed as $\widehat{H}=\frac{1}{2\ln 2}\bigl(\ln\bar{A}-\ln\bar{B}\bigr),$ and its derivatives with respect to $\bar{U}_{n,N}$ are given by
$\frac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}$ $=\frac{h^{2}}{2\ln 2}\Bigl(\frac{1}{\bar{B}}-\frac{4}{\bar{A}}\Bigr),$ $\frac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}$ $=0,$
$\frac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}$ $=-\frac{1}{2\ln 2}\,\frac{1}{\bar{B}},$ $\frac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}$ $=\frac{1}{2\ln 2}\,\frac{1}{\bar{A}}.$
@@
Evaluating at $U_{\infty}$ and using $A=2^{2H}B$ and $B=\gamma^{2}D_{H}$ with $D_{H}:=h^{2H}\bigl(2^{2H-1}-1\bigr)$, yields
$\frac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}$ $=\frac{h^{2}}{2\ln 2}\,\frac{2^{2H}-4}{2^{2H}\gamma^{2}D_{H}},\quad\frac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}=0,$
$\frac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}$ $=-\frac{1}{2\ln 2}\,\frac{1}{B},\quad\frac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}=\frac{1}{2\ln 2}\,\frac{1}{2^{2H}B}.$
@@
Next, for the estimator $\widehat{\gamma}^{2}=\dfrac{\bar{B}}{D_{\widehat{H}}}$, the chain rule gives for any component $u\in\{\bar{V}_{n,N},\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N}\}$,
$\frac{\partial\widehat{\gamma}^{2}}{\partial u}=\frac{\partial_{u}\bar{B}}{D_{\widehat{H}}}-\frac{\bar{B}\,D_{\widehat{H}}^{\prime}}{D_{\widehat{H}}^{2}}\,\frac{\partial\widehat{H}}{\partial u},$
@@
where $D_{H}^{\prime}:=\frac{\partial D_{H}}{\partial H}=h^{2H}\Bigl[2\ln(h)\bigl(2^{2H-1}-1\bigr)+(2\ln 2)\,2^{2H-1}\Bigr].$ Using $\dfrac{\partial\bar{B}}{\partial\bar{V}_{n,N}}=-h^{2}$, $\dfrac{\partial\bar{B}}{\partial\bar{\eta}_{n,N}}=1$, and $\dfrac{\partial\bar{B}}{\partial\bar{\xi}_{n,N}}=\dfrac{\partial\bar{B}}{\partial\bar{\zeta}_{n,N}}=0$, and evaluating at $U_{\infty}$ with $B=\gamma^{2}D_{H}$ and $A=2^{2H}B$, we obtain
@@
$\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}$ $=-\frac{h^{2}}{D_{H}}+\frac{h^{2}}{2\ln 2}\Bigl(\frac{4}{2^{2H}}-1\Bigr)\frac{D_{H}^{\prime}}{D_{H}^{2}},$
$\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{\xi}_{n,N}}\Big|_{U_{\infty}}$ $=0,$
$\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}$ $=\frac{1}{D_{H}}+\frac{D_{H}^{\prime}}{2\ln 2}\,\frac{1}{D_{H}^{2}},$
$\frac{\partial\widehat{\gamma}^{2}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}$ $=-\frac{D_{H}^{\prime}}{2\ln 2}\,\frac{1}{2^{2H}D_{H}^{2}}.$
Finally, for $\widehat{\sigma}^{2}=\dfrac{\bar{\xi}-h^{2}\bar{V}}{h}-\dfrac{\widehat{\gamma}^{2}h^{2\widehat{H}}}{h}$, we use
$\frac{\partial}{\partial u}\bigl(\widehat{\gamma}^{2}h^{2\widehat{H}}\bigr)=\Bigl(\frac{\partial}{\partial u}\widehat{\gamma}^{2}\Bigr)h^{2\widehat{H}}+\widehat{\gamma}^{2}h^{2\widehat{H}}(2\ln h)\frac{\partial\widehat{H}}{\partial u}.$
@@
Evaluating at $U_{\infty}$, which means that $h^{2\widehat{H}}\to h^{2H}$ and $\gamma^{2}h^{2H}=\dfrac{B}{D_{H}}$, we obtain
@@
$\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{V}_{n,N}}\Big|_{U_{\infty}}$ $=-h+h^{2H-1}\Bigl[\frac{h^{2}}{D_{H}}-\frac{h^{2}}{2\ln 2}\Bigl(1-\frac{4}{2^{2H}}\Bigr)\Bigl(-\frac{D_{H}^{\prime}}{D_{H}^{2}}+\frac{2\ln h}{D_{H}}\Bigr)\Bigr],$
@@
$\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{\eta}_{n,N}}\Big|_{U_{\infty}}$ $=-h^{2H-1}\frac{1}{D_{H}}\Bigl[1-\frac{2\ln h}{2\ln 2}+\frac{D_{H}^{\prime}}{2\ln 2}\,\frac{1}{D_{H}}\Bigr],$
$\frac{\partial\widehat{\sigma}^{2}}{\partial\bar{\zeta}_{n,N}}\Big|_{U_{\infty}}$ $=h^{2H-1}2^{-2H}\Bigl[\frac{D_{H}^{\prime}}{2\ln 2}\,\frac{1}{D_{H}^{2}}-\frac{2\ln h}{2\ln 2}\,\frac{1}{D_{H}}\Bigr].$